Circular No. 5657

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 R. H. McNaught, Anglo-Australian Observatory, reports the following

measurement obtained from an Uppsala Southern Schmidt telescope film obtained on Nov. 13.6 UT: α = 5^{h}19^{m}36^{s}.88, δ = −68^{o}57′31′′.4 (equinox 1950.0, uncertainty 0′′.3 in α, 0′′.2 in δ). A nearby star of mag 13 has end figures 24^{s}.16, 34′′.1.

S. N. Shore, Goddard High Resolution Spectrograph team and Computer Sciences Corporation; S. Starrfield, Arizona State University; R. Gonzalez-Riestra, International Ultraviolet Explorer (IUE) Observatory, European Space Agency, Madrid; and G. Sonneborn, Goddard Space Flight Center, NASA, report: “Low-dispersion spectrograms (115.0-330.0 nm, resolution 0.6 nm) of this object were obtained with the IUE satellite on Nov. 13.3 and 19.2 UT. The spectra are all characteristic of an optically thick shell with a relatively slow expansion velocity. The Nov. 13 spectra were probably taken at ultraviolet minimum; the Fine Error Sensor magnitudes were 10.4 on Nov. 13.2 and 11.9 on Nov. 19.1. The integrated fluxes in units of 10^{−11} erg cm^{−2} s^{−1} for 120-200 nm were: Nov. 13.4, 0.34; Nov. 19.3, 3.57; and for 210-330 nm were: Nov. 13.2, 10.9; Nov. 19.2, 11.9. The ultraviolet spectrum showed a continuum sharply attenuated shortward of 170 nm by absorption from an optically thick shell. The absorption is as severe as is typically seen in the earliest ultraviolet spectra of classical novae, but the optical depth is decreasing quickly and continuum windows near 150 and 180 nm have appeared. The nova is evolving at roughly constant bolometric luminosity which, for an assumed distance of 55 kpc and E(B−V) = 0.15, gives a luminosity of ∼ 30 000 L_{⊙} on Nov. 19.2. Mg II was in emission, with FWHM = 1800 km/s; Al II 264.0-nm is also present. No other prominent single emission lines were present in these early spectra. Ultraviolet observations are continuing.”

NOVA SAGITTARII 1992 No. 3 Visual and photovisual magnitude estimates (cf. IAUC 5652): Nov.

7.10 UT, 9.5 (P. Collins, Boulder, CO); 8.21, 9.5 (Collins); 14.71, 10.5 (A. Boattini, Piazzano, Italy); 15.06, 10.21 (W. Liller, Viña del Mar, Chile; CCD); 16.07, 10.34 (Liller; CCD).

1992 November 19 Daniel W. E. Green